# Robust Physics-Informed Neural Network Approach for Estimating Heterogeneous Elastic Properties from Noisy Displacement Data

### v1

@@ 1 Introduction
Quantifying the spatial distributions of elastic properties, specifically Young’s modulus and Poisson’s ratio is crucial in numerous applications, including biomedical imaging. Young’s modulus characterizes a material’s local resistance to elastic (reversible) axial deformation, while Poisson’s ratio quantifies the coupling between axial and transverse deformations. When external loads are applied, an internal stress field is established within the solid material, resulting in deformation (i.e., displacement) patterns that depend directly on spatial distributions of the elastic properties. Accurate spatial characterization of elastic properties enables precise disease diagnosis (e.g., cardiovascular and airway diseases[1, 2, 3, 4, 5], cancer.[6, 7, 8]), development biomedical devices and tissue engineering scaffolds,[9, 10, 11] assessment of structural integrity (e.g., bone health,[12] engineered and additive manufactured parts.[13, 14, 15]), and precise computational mechanics modeling.[16, 17] Typically, elastic properties vary spatially within these materials, necessitating advanced methods to characterize their heterogeneity.
@@ 4.4 Neural Network Architecture
Figure S4: Young’s modulus ($E$) and Poisson’s ratio $(\nu)$ prediction and error map of different models at a signal-to-noise ratio of 500. The models were trained using the displacement data with a higher noise level: SNR 500. (i) IE-PINN (Proposed): Model that fits in both displacement and strain. (ii) Model with fitting only displacement data without extra strain fitting and (iii) Elastnet: Model without fitting in both displacement and strain data[64]. When the noise increases, relying solely on fitting displacement is less effective than additional fitting strain.
@@
The datasets for simulation are generated from the published data in the work by Chen and Gu[46, 64]. The noise is imputed to the raw displacement data as the observed noisy data based on the formulation below.
@@
The loading boundary condition refers to the applied force that can cause deformation in an object. This training dataset is generated using a finite element model of an elastic quadrilateral body, where its elasticity is defined by one of 12 zodiac shapes in Young’s modulus ($E$), and Poisson’s ratio $(\nu)$[46]. The top and bottom boundaries of the object are free. In contrast, the left boundary is fixed to restrict horizontal movement (in the x-direction) to zero, allowing only horizontal movement in the right boundary and vertical movement (in the y-direction). The objective of the simulation is to achieve a displacement of 1% of the body length at the right boundary, which is expected to result in an average normal strain along the x-direction $(\varepsilon_{xx})$ of 1%. Under these conditions, the applied force that defines the loading boundary condition can generally be obtained by the total traction force derived from the stress (calculated from the true boundary strain and elasticity). Since the goal is to produce horizontal displacement at the right boundary, the boundary stress in the x-direction is used to obtain the applied force, as shown in the following equation:

### v2

@@ 1 Introduction
Quantifying the spatial distributions of elastic properties, specifically Young’s modulus and Poisson’s ratio is crucial in numerous applications, including biomedical imaging. Young’s modulus characterizes a material’s local resistance to elastic (reversible) axial deformation, while Poisson’s ratio quantifies the coupling between axial and transverse deformations. When external loads are applied, an internal stress field is established within the solid material, resulting in deformation (i.e., displacement) patterns that depend directly on spatial distributions of the elastic properties. Accurate spatial characterization of elastic properties enables precise disease diagnosis (e.g., cardiovascular and airway diseases,[1, 2, 3, 4, 5] cancer.[6, 7, 8]), development biomedical devices and tissue engineering scaffolds,[9, 10, 11] assessment of structural integrity (e.g., bone health,[12] engineered and additive manufactured parts.[13, 14, 15]), and precise computational mechanics modeling.[16, 17] Typically, elastic properties vary spatially within these materials, necessitating advanced methods to characterize their heterogeneity.
@@ 4.4 Neural Network Architecture
Figure S4: Young’s modulus ($E$) and Poisson’s ratio $(\nu)$ prediction and error map of different models at a signal-to-noise ratio of 500. The models were trained using the displacement data with a higher noise level: SNR 500. (i) IE-PINN (Proposed): Model that fits in both displacement and strain. (ii) Model with fitting only displacement data without extra strain fitting and (iii) Elastnet: Model without fitting in both displacement and strain data.[64] When the noise increases, relying solely on fitting displacement is less effective than additional fitting strain.
@@
The datasets for simulation are generated from the published data in the work by Chen and Gu.[46, 64] The noise is imputed to the raw displacement data as the observed noisy data based on the formulation below.
@@
The loading boundary condition refers to the applied force that can cause deformation in an object. This training dataset is generated using a finite element model of an elastic quadrilateral body, where its elasticity is defined by one of 12 zodiac shapes in Young’s modulus ($E$), and Poisson’s ratio $(\nu)$.[46] The top and bottom boundaries of the object are free. In contrast, the left boundary is fixed to restrict horizontal movement (in the x-direction) to zero, allowing only horizontal movement in the right boundary and vertical movement (in the y-direction). The objective of the simulation is to achieve a displacement of 1% of the body length at the right boundary, which is expected to result in an average normal strain along the x-direction $(\varepsilon_{xx})$ of 1%. Under these conditions, the applied force that defines the loading boundary condition can generally be obtained by the total traction force derived from the stress (calculated from the true boundary strain and elasticity). Since the goal is to produce horizontal displacement at the right boundary, the boundary stress in the x-direction is used to obtain the applied force, as shown in the following equation: